OpenStudy (geekfromthefutur):

Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours. Use complete sentences

OpenStudy (geekfromthefutur):

@jdoe0001 @amistre64 @Luigi0210

OpenStudy (geekfromthefutur):

@ganeshie8

OpenStudy (anonymous):

can i help

OpenStudy (geekfromthefutur):

Sure i been up for 4 hours stuck on this question.

OpenStudy (anonymous):

really

OpenStudy (geekfromthefutur):

yup well altogther today 2 hours yesterday 2 aswell so equal up to 4

OpenStudy (anonymous):

Bullet Train = 320 km / hour TGV train = 89.44 meters / second = 89.44 X 3600 meters / hour = 89.44 X 3600 /1000 km/ hour = 89.44 X 3.6 = 321.984 km/ hour TGV train is faster Ellen's train has to go faster than 320 km/ hour but less than 321.984 km/ hour i.e around 321 km/ hour to be the second fastest train ANSWER Distance in 2.5 hours = 321 X 2.5 = 802.5 km ANSWER

OpenStudy (geekfromthefutur):

really yahoo answers

OpenStudy (anonymous):

come on that is easier dude

OpenStudy (geekfromthefutur):

Well how am i suppose to put that in complete sentences??

OpenStudy (geekfromthefutur):

nvm

OpenStudy (anonymous):

okay atleast i gave u and answer

OpenStudy (geekfromthefutur):

Btw this is 9th grade ur prob in 6th or 7th grade

OpenStudy (anonymous):

i am in 8 th

OpenStudy (geekfromthefutur):

@phi @whpalmer4 @mathslover

OpenStudy (phi):

what part of the problem confuses you? you are given the speed of the TGV train is 89.44 meters per second. convert that to km/hour

OpenStudy (geekfromthefutur):

what did i had to convert and wanted to know if i had to multiply or divide

OpenStudy (phi):

to convert units, you start by writing down that you have \[ \frac{meters}{second}\] you want \[ \frac{km}{hour} \] you need "conversion factors" to multiply the first fraction by to get the second fraction if we look just at the units, to change meters to km we want to multiply by \[ meters \cdot \frac{km}{meters} \] the idea is that the meters "cancel" so you need a factor of km per meter if you know 1000 meters per km or \( 1000 \frac{meters}{km}\) that is "upside down". if we flip it we get \[ \frac{1}{1000} \frac{km}{meter} \]

OpenStudy (phi):

does that part make any sense?

OpenStudy (geekfromthefutur):

Now i se how you are at 99 lol yea it make alot of sense

OpenStudy (phi):

so far we have \[ 89.44 \frac{\cancel{meters}}{sec} \cdot \frac{1}{1000} \frac{km}{\cancel{meters}}= \frac{89.44}{1000} \frac{km}{sec} \]

OpenStudy (geekfromthefutur):

So its 89.44 seconds x 1/1000km=89.44/1000 km/sec

OpenStudy (phi):

89.44 meters/second * 1/1000 km/meter the meters cancel (divide out) and leave km/sec

OpenStudy (geekfromthefutur):

O ok

OpenStudy (phi):

we want to change seconds to hours 60 seconds = 1 minute 60 minute = 1 hour 60*60 seconds = 1 hour 3600 seconds = 1 hour we have 1 hour per 3600 seconds \[\frac{1 \text{ hour}}{3600 \text{ sec}} = \frac{1}{3600} \frac{hour}{sec}\]

OpenStudy (geekfromthefutur):

Oh ok

OpenStudy (phi):

your problem is (so far) \[ \frac{89.44}{1000} \frac{km}{sec} \] if you want to "cancel" the secs what should we multiply by ? hour/sec or sec/hour ?

OpenStudy (geekfromthefutur):

sec/hour?

OpenStudy (phi):

yes, because sec "up top" will cancel the sec "down below" we know the conversion factor \[ \frac{1}{3600} \frac{hour}{sec} \] but that is upside down. Flip it to \[ 3600 \frac{sec}{hour} \] multiply your problem by that factor

OpenStudy (phi):

btw 1/3600 flipped is 3600/1 but when we divide by 1, we don't bother to write it. we just write 3600

OpenStudy (geekfromthefutur):

So we have 3600 is how far the train can go in 2.5 seconds?

OpenStudy (phi):

no. what we know is there is 3600 seconds per hour the "conversion factor" for seconds/ hour is 3600 we use the conversion factor. \[ \frac{89.44}{1000} \frac{km}{sec} \cdot 3600 \frac{sec}{hour} \]

OpenStudy (geekfromthefutur):

Ok so distance = speed • time the distance would be not yet found But the spped is about to be but the time is 89.44/1000 km/sec * sec/hour

OpenStudy (phi):

But the spped is about to be but the time is 89.44/1000 km/sec * sec/hour can you say that more clearly ?

OpenStudy (geekfromthefutur):

Ok what im saying that distance is not yet solved but the speed is about to turn into time

OpenStudy (phi):

we have been working on only one thing: change speed measured in meters per second to speed measured in kilometers per hour we will still have speed, but different units. the final step in this "conversion" is to simplify this expression \[ \frac{89.44}{1000} \frac{km}{sec} \cdot 3600 \frac{sec}{hour} \]

OpenStudy (geekfromthefutur):

O ok

OpenStudy (phi):

notice you don't just multiply by sec/hour. you need to multiply by a conversion factor (number) that has units of sec/hour. we know that 3600 sec per hour or 3600 sec/hour is the factor

OpenStudy (geekfromthefutur):

Ok i understand that now

OpenStudy (phi):

what do you get after simplifying \[ \frac{89.44}{1000} \frac{km}{sec} \cdot 3600 \frac{sec}{hour} \]?

OpenStudy (geekfromthefutur):

268.32

OpenStudy (phi):

your calculator is broken. try again

OpenStudy (geekfromthefutur):

lol ok

OpenStudy (geekfromthefutur):

321.984??

OpenStudy (phi):

yes. but with these problems *always* include the units can you do that ?

OpenStudy (geekfromthefutur):

I can try nothing like putting effort in anything

OpenStudy (phi):

\[ \frac{89.44}{1000} \frac{km}{sec} \cdot 3600 \frac{sec}{hour} = 321.984 \frac{km}{sec} \cdot \frac{sec}{hour} \] simplify the units (by noticing that sec/sec = 1)

OpenStudy (geekfromthefutur):

O ok

OpenStudy (geekfromthefutur):

What about km is that what i have to figure out?

OpenStudy (phi):

first, simplify \[ 321.984 \frac{km}{sec} \cdot \frac{sec}{hour} \]

OpenStudy (geekfromthefutur):

sec/hour= 1 km/sec=

OpenStudy (phi):

Here is the rule: if you have the *same thing* in the top and bottom, you can erase both

OpenStudy (geekfromthefutur):

ok

OpenStudy (phi):

do you see "sec" is up top and down below? that means you can erase both of them what is left over ?

OpenStudy (geekfromthefutur):

Km/sec

OpenStudy (phi):

no. you erase sec from the top and you erase sec from the bottom what is left in\[ 321.984 \frac{km}{sec} \cdot \frac{sec}{hour} \]

OpenStudy (geekfromthefutur):

*Km/hour

OpenStudy (phi):

yes, we have the speed is 321.984 km/hr

OpenStudy (geekfromthefutur):

Wow

OpenStudy (phi):

now for the rest of the question TGV train speed is 321.984 km/hr Shinkansen Bullet Train 320 km/hr you want a train that is second fastest (slower than TGV and faster than Shinkansen)

OpenStudy (phi):

that means pick a speed between those two numbers

OpenStudy (geekfromthefutur):

Ok

OpenStudy (geekfromthefutur):

TGV train speed is 321.984 km/hr

OpenStudy (phi):

yes but we need to answer what speed her new train must go to be the second fastest in this group.

OpenStudy (phi):

Ellen is building a new train. She wants it to be slower than TGV but faster than the other

OpenStudy (geekfromthefutur):

It need to go 321.000

OpenStudy (phi):

yes, that is the easiest speed to pick. but it could go lots of different speeds: 321.1 or 321.123 or 320.5 or.... There are an infinite number of speeds it could be. but 321 is definitely the "nicest" number we can pick.

OpenStudy (phi):

Last part of the question distance = speed • time in 2.5 hours at speed = 321 km/hour

OpenStudy (phi):

you should put the numbers and the units into the equation (be sure to include the units)

OpenStudy (geekfromthefutur):

Distance= speed * time = 321.00km/hour

OpenStudy (phi):

you need to include * time time is 2.5 hours

OpenStudy (geekfromthefutur):

oh yea i meant to add that i forgot i knew i forgot something Distance= speed x time =321.00km/hour x 2.5hours

OpenStudy (phi):

ok, that looks good. you have \[ Distance = 321.00 \frac{km}{hour} \cdot 2.5 \text{ hour} \] or \[ Distance = 321.00 \cdot 2.5 \frac{km}{hour} \cdot \text{ hour} \] simplify the units Here is the rule: if you have the *same thing* in the top and bottom, you can erase both

OpenStudy (phi):

remember that hour is the same as hour/1 so the "hour" is up top.

OpenStudy (geekfromthefutur):

The first 1.

OpenStudy (phi):

\[ Distance = 321.00 \cdot 2.5 \frac{km}{hour} \cdot \frac{ hour} {1}\] simplify the units Here is the rule: if you have the *same thing* in the top and bottom, you can erase both

OpenStudy (geekfromthefutur):

804.960

OpenStudy (phi):

First simplify the units

OpenStudy (geekfromthefutur):

km/hour km/1

OpenStudy (geekfromthefutur):

i did with the 1/km cuz i took out the hour and turned it into 1

OpenStudy (phi):

Here is the rule: if you have the *same thing* in the top and bottom, you can erase both \[ Distance = 321.00 \cdot 2.5 \frac{km}{\cancel{hour}} \cdot \frac{\cancel{\text{ hour}}}{1} \] like that. we don't need to show the 1 (divide by 1 is not usually written down) so we get \[ Distance = 321.00 \cdot 2.5\ km \]

OpenStudy (geekfromthefutur):

Oh yea my bad i was going to erase the 1 and hour sense thats a part of an hour now i understand that

OpenStudy (phi):

now multiply 321 * 2.5 be careful.

OpenStudy (geekfromthefutur):

80.25

OpenStudy (geekfromthefutur):

????

OpenStudy (phi):

no. First, always write down the units (the km) or you might forget second, 321 * 2.5 will be bigger than 321 (so it can't be 80 !)

OpenStudy (geekfromthefutur):

oh ok

OpenStudy (geekfromthefutur):

802.5?

OpenStudy (phi):

yes, but with the units

OpenStudy (geekfromthefutur):

802.5km

OpenStudy (phi):

yes.

OpenStudy (phi):

to finish this Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. the answer is change the speed in m/sec to km/hr You can look up above to see how we did that. then compare the speed of the two trains and pick a speed in between the two trains' speeds finally, show the work to find the distance traveled by Ellen's train going at 321 km/hr for 2.5 hours

OpenStudy (geekfromthefutur):

Wait wat that a little confusing @phi

OpenStudy (geekfromthefutur):

Ellen's train must go 802.5km

OpenStudy (geekfromthefutur):

to be the second fastest

OpenStudy (phi):

Ellen's train goes 802.5 km in 2.5 hours. That is a distance. But the question first asks Explain to Ellen how to find what speed speed is km/hour (speed is how fast you move)

OpenStudy (geekfromthefutur):

ok